r11007

Hobart 26m

Disk VSN: GSFC+011 | Data volume at beginning: 5.77 TB

  • 1700UT Experiment started okay (Lim)
  • 2306UT Schedule halted for cryo repair (JMc)
  • 0038UT SChedule resumed. Cryognics active, still cooling but should reach the target level in a few hors. Pressure ok (JMc)